# biblio.bib

### v1

@@ 3. Proof of the upper bound in Theorem 2
An argument of Lieb shows that $|x_{j}-x_{k}|\geqslant c(m_{*}/\rho)^{1/3}$ for some constant $c>0$ independent of $N$ and all $j\neq k$ (see [Lewin-22, Lemma 25]). It will be useful to consider in addition a different configuration of points $Y=\{y_{1},...,y_{N}\}\subset C_{\ell}$ which also satisfies $\sum_{j=1}^{N}y_{j}=0$ and the additional condition that $\mathrm{d}(y_{j},\partial C_{\ell})\geqslant c(m_{*}/\rho)^{1/3}$, that is, staying far away from the boundary. For concreteness we can place them on a lattice of side length $\varepsilon(m_{*}/\rho)^{1/3}$ for some $\varepsilon>0$ small enough. We can do this in such a way that $|y_{j}-y_{k}|\geqslant c(m_{*}/\rho)^{1/3}$ for all $j\neq k$.
Now we give ourselves an optimal set $\Omega_{*}\subset B(0,r_{*})$ for $\mu_{*}$ in (9), of volume $|\Omega_{*}|=m_{*}$. We then consider the points
@@
of side length $(2k+3)\ell$ and the total number of points is $K:=\#Z=(2k+3)^{3}N$. By construction the points of $Z$ are at distance $\geqslant c(m_{*}/\rho)^{1/3}$ from each other and from the boundary of $\Lambda$. We used the external layer to ensure that the points are all strictly inside $\Lambda$. Next we consider the trial set
@@
Next we average this bound over the independent rotations $R_{z}$’s. By Newton’s theorem, we have
$\fint_{SO(3)}\fint_{SO(3)}D({\mathds{1}}_{z+R\Omega_{*}},{\mathds{1}}_{z^{\prime}+R^{\prime}\Omega_{*}})\,\mathrm{d}R\,\mathrm{d}R^{\prime}=\frac{m_{*}^{2}}{|z-z^{\prime}|}$
@@
$\qquad\qquad=m_{*}\int_{\Lambda}\frac{\mathrm{d}y}{|z-y|}+\fint_{SO(3)}\int_{R\Omega_{*}}\int_{B(0,r_{*})}\left(\frac{1}{|x-y|}-\frac{1}{|y|}\right)\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}R$
@@
Indeed, if $\xi$ is not a neighbor of $\xi_{0}$ this is [AnaLew-20, Lemma 2.1] and the constant $C_{N}$ can be bounded by $CN^{2}$. If the two cubes touch, we need to use that $Y$ is at distance $(m_{*}/\rho)^{1/3}=\ell N^{-1/3}$ to the boundary and we obtain the bound
@@
because we have the uniform bound ${\mathds{1}}_{C_{\ell}}\ast{|x|}^{-1}\leqslant C\ell^{2}$. The latter follows for instance from Newton’s theorem, using
@@
is the energy of one cube from the outer layer, and $(2k+3)^{3}-(2k+1)^{3}$ equals the number of such cubes. Next we divide by the volume $|\Lambda|=(2k+3)^{3}\ell^{3}=(2k+3)^{3}m_{*}N/\rho$ and we take the limit $k\to\infty$ with all the other parameters fixed. For the Jellium term we use the convergence to the periodic energy as stated for instance in [Lewin-22, Lemma 32]. We obtain
@@
where $M_{\mathbb{Z}^{3}}=\lim_{x\to 0}(G_{1}(x)-|x|^{-1})$ is the Madelung constant of the lattice $\mathbb{Z}^{3}$. As a last step we take the limit $\ell\to\infty$, using that the periodic Jellium problem converges to $e_{\rm Jel}$ in (16), by [LewLieSei-19b, Section V] and [Lewin-22, Theorem 40]. There is an additional factor $(\rho/m_{*})^{4/3}$ since the effective density of the periodic problem is $\rho/m_{*}$. Our final bound is
@@ 4. Proof of the lower bound in Theorem 2
The name “grand-canonical” comes from statistical physics and $\mu_{*}$ is interpreted as a “chemical potential”.
@@
In this first step we localize our problem to sets of finite size. For this we use the Graf–Schenker inequality [GraSch-95, HaiLewSol_thermo-09]. Locally we get the grand-canonical problem introduced in (26) because although we start with a globally neutral system ($|\Omega|=\rho|\Lambda|$), neutrality does not hold locally. The precise statement is the following.
@@
We show in the next lemma that there is a constant $C$ such that
@@
Next we take a large cube $\Lambda_{L}:=(-L/2,L/2)^{3}$ and any domain $\Omega\subset\Lambda_{L}$. The Graf–Schenker inequality from [GraSch-95] states that
@@
For fixed $g$ we expand the integrand in the remaining double integral. Together with (28) and (29), we obtain for all $\Omega\subset\Lambda_{L}$
@@
The right side is the Jellium energy for particles with a charge $m_{*}$, in a fixed tetrahedron of volume $A^{3}$ and a uniform background of density 1. It is grand-canonical because we optimize over the number of particles $n$ instead of improsing the neutrality condition $m_{*}n=A^{3}$. It is however known that the optimal $n$ satisfies $n=A^{3}/m_{*}+o(A^{2})$. Fluctuations about the neutral case are negligible compared to the surface, otherwise the resulting energy is modified to the order of the volume (see [LieLeb-72, LieNar-75, GraSch-95b] and [Lewin-22, Cor. 50]).
We call (33) the “ultra-dilute” limit because we expect that the solution to the left side involves finitely many droplets (of the order of $A^{3}$) located at distances of order $\rho^{-1/3}$ in the very large domain $\ell\Delta$ of side length $\ell=A\rho^{-1/3}$. A similar limit was considered for the canonical problem in [ChoPel-10, KnuMurNov-16, EmmFraKon-20]. After rescaling everything by $\rho^{1/3}$ the droplets look like points and we find the right side.
@@
For $\rho=0$ we just recover the result from [FraKilNam-16], of which the lemma is an easy extension. When $\rho>0$ the inequality (34) says that $|\Omega|$ is at most of the same order as the background charge $\rho\,|\Lambda|\approx\rho\ {\rm diam}(\Lambda)^{3}$.
@@
Our proof follows the method in [FraKilNam-16]. For any $\nu\in\mathbb{S}^{2}$ and $t\in\mathbb{R}$, we define the half spaces and the plane
@@
Upon translating everything we can assume that $\Lambda\subset B(0,R)$ with $R=\text{diam}(\Lambda)$ (in fact, by Jung’s theorem [Federer-96, Thm. 2.10.41] we could take $R=\text{diam}(\Lambda)\sqrt{3/8}$ which would slightly improve the constant $16\pi$ in the statement of the lemma). Then we know that $\mathcal{H}^{\pm}_{\nu,\pm t}\cap\Lambda=\emptyset$ for all ±$\pm t\geqslant R$. For every $t\in(-R,R)$, we split $\Omega$ into
@@
by Newton’s theorem as in (23). Next we integrate over $\nu\in\mathbb{S}^{2}$ and $t\in(-R,R)$, which gives as in [FraKilNam-16]
@@
Writing $A^{3}/m_{*}=N+t=(1-t)N+t(N+1)$ with $N\in\mathbb{N}_{0}$ and $10\leqslant t<$ 1 we can then bound the first minimum by the convex combination
@@
Next we remark as in [DauLie-83] that the right side is linear with respect to each $m_{*}^{(j)}$, when all the other ones are fixed. We can thus replace all the $m_{*}^{(j)}$ by either 00 or the maximal value $m_{*}$ and decrease the right side. When they are replaced by 00 this is equivalent to discarding the corresponding points $y_{j}$, that is, to decrease $K$. Relabelling the points $y_{j}$ so that the $K^{\prime}$ first end up having mass $m_{*}$, we find that
@@
as we wanted. The grand-canonical Jellium problem has the same thermodynamic limit as its canonical counterpart in (16) by [Lewin-22, Thm. 24]. This concludes the proof of Theorem 2.∎
@@ Appendix A Proof of Theorem 1 on the thermodynamic limit
In this appendix we describe how to prove the existence of the thermodynamic limit stated in (6), using the method introduced by Lieb, Lebowitz and Narnhofer in [LieLeb-72, LieNar-75]. It is also presented in [LieSei-09, Chap. 14].
The proof is, as usual, divided into three steps. The first step consists of establishing the existence of the limit for a special sequence of balls $B_{j}$ growing exponentially fast. This is done by packing any $B_{K}$ by a certain disjoint union of copies of the smaller balls $B_{0},...,B_{K-1}$ in such a way that the remaining space has a volume negligible compared with the large volume $|B_{K}|$. The existence of such packings was proved first in [LieLeb-72] and has become known as a “swiss cheese theorem”. In the next two steps we use this special sequence to obtain the limit for any sequence of sets $\Lambda_{n}\nearrow\mathbb{R}^{3}$, satisfying suitable regularity conditions. Upper and lower bounds are proved using a similar swiss cheese for the inside or the outside of $\Lambda_{n}$.
@@ A.1. Construction of a quadrupole layer
We start by presenting a technical result that will be used in each of the three steps. It will allows us to deal with the neighborhood of the boundary of the sets we will manipulate. Recall that for the slowly-decaying Coulomb potential, a small charge disturbance can have a large effect far away. When constructing trial sets, it is therefore important to ensure that the charges locally balance well. Local neutrality is not enough and in the following proposition we will achieve an additional vanishing dipole condition. Handling this situation for the liquid drop model is harder than what was previously done in [LieLeb-72, LieNar-75] and this is our main contribution in this appendix.
@@
Let $\rho\leqslant 1/2$. Let $\Lambda\subset\mathbb{R}^{3}$ be an $(r,L)$–Lipschitz open set with compact boundary. Let $C_{z}=\varepsilon z+(-\varepsilon/2,\varepsilon/2)^{3}$ be a tiling of cubes of size $\varepsilon>0$, with $z\in\mathbb{Z}^{3}$. Define the outer cubic boundary layer of $\Lambda$ by
@@
Recall that an open set $\Lambda\subset\mathbb{R}^{3}$ is $(r,L)$–Lipschitz if for any $x\in\partial\Lambda$, the set $\Lambda\cap B(x,r)$ is, in an appropriate coordinate system, equal to the epigraph of an $L$–Lipschitz function. The interpretation of the condition (36) is that the charge distribution ${\mathds{1}}_{\Omega_{\alpha}}-\rho{\mathds{1}}_{\Lambda_{\alpha}}$ has no charge and no dipole moment. It can be thought as a small quadrupole, hence the title of the proposition. Recall that in our situation the quadrupole moment is the $3\times 3$ symmetric trace-less matrix
@@
Finally we come to the most interesting type of $\Lambda_{\alpha}$. We define each remaining $\Lambda_{\alpha}$ as the union of one $C_{z}$ with $z\in\mathcal{Z}_{i}$ together with all the small partial cubes $C^{\prime}_{z^{\prime}}\setminus\Lambda$ associated to it. The picture is therefore that we have a large full cube of volume $\varepsilon^{3}$ together with several small partial cubes of volume $\leqslant\varepsilon^{3}/K^{3}$ (Figure 2). The $(r,L)$–Lipschitz property implies the following two facts, for $\varepsilon\leqslant r/C$:
@@
The constant $C$ only depends on $L$. The first claim (a) follows for instance from the fact that at any point $x\in\partial\Lambda$ there is a cone of height comparable with $r$ and opening angle bounded from below in terms of $L$, that is completely contained in $\mathbb{R}^{3}\setminus\Lambda$ (see for instance [Grisvard, Thm. 1.2.2.2]). Hence for $\varepsilon\leqslant r/C$ this cone will contain a full cube $C_{z}$. For the second claim (b), we use that the $\eta$–thickened boundary inside a ball of radius $r^{\prime}$ satisfies

### v2

@@ 3. Proof of the upper bound in Theorem 2
An argument of Lieb shows that $|x_{j}-x_{k}|\geqslant c(m_{*}/\rho)^{1/3}$ for some constant $c>0$ independent of $N$ and all $j\neq k$ (see [Lew22, Lemma 25]). It will be useful to consider in addition a different configuration of points $Y=\{y_{1},...,y_{N}\}\subset C_{\ell}$ which also satisfies $\sum_{j=1}^{N}y_{j}=0$ and the additional condition that $\mathrm{d}(y_{j},\partial C_{\ell})\geqslant c(m_{*}/\rho)^{1/3}$, that is, staying far away from the boundary. For concreteness we can place them on a lattice of side length $\varepsilon(m_{*}/\rho)^{1/3}$ for some $\varepsilon>0$ small enough. We can do this in such a way that $|y_{j}-y_{k}|\geqslant c(m_{*}/\rho)^{1/3}$ for all $j\neq k$.
We consider the points
@@
of side length $(2k+3)\ell$ and the total number of points is $K:=\#Z=(2k+3)^{3}N$. By construction the points of $Z$ are at distance $\geqslant c(m_{*}/\rho)^{1/3}$ from each other and from the boundary of $\Lambda$. We used the external layer to ensure that the points are all strictly inside $\Lambda$.
Now we give ourselves an optimal set $\Omega_{*}\subset B(0,r_{*})$ for $\mu_{*}$ in (9), of volume $|\Omega_{*}|=m_{*}$, and consider the trial set
@@
Next we average this bound over the independent rotations $R_{z}$’s. Recall that Newton’s theorem [LL01, Thm. 9.7] implies
$\fint_{SO(3)}\int_{\mathbb{R}^{d}}\frac{{\mathds{1}}_{z+R\Omega_{*}}(y)}{|x-y|}\,\mathrm{d}y\,\mathrm{d}R=\frac{m_{*}}{|x-z|}\qquad\text{for $|x-z|\geqslant r_{*}$.}$ (23)
Hence
$\fint_{SO(3)}\fint_{SO(3)}D({\mathds{1}}_{z+R\Omega_{*}},{\mathds{1}}_{z^{\prime}+R^{\prime}\Omega_{*}})\,\mathrm{d}R\,\mathrm{d}R^{\prime}=\frac{m_{*}^{2}}{2|z-z^{\prime}|}$
@@
$\qquad\qquad=m_{*}\int_{\Lambda}\frac{\mathrm{d}y}{|z-y|}+\fint_{SO(3)}\int_{R\Omega_{*}}\int_{B(0,r_{*})}\left(\frac{1}{|x-y|}-\frac{1}{|y|}\right)\mathrm{d}y\,\mathrm{d}x\,\mathrm{d}R$
@@
Indeed, if $\xi$ is not a neighbor of $\xi_{0}$ this is [AL20, Lemma 2.1] and the constant $C_{N}$ can be bounded by $CN^{2}$. If the two cubes touch, we need to use that $Y$ is at distance $(m_{*}/\rho)^{1/3}=\ell N^{-1/3}$ to the boundary and we obtain the bound
@@
because we have the uniform bound ${\mathds{1}}_{C_{\ell}}\ast{|x|}^{-1}\leqslant C\ell^{2}$. The latter follows for instance from Newton’s theorem [LL01, Thm. 9.7], using
@@
is the energy of one cube from the outer layer, and $(2k+3)^{3}-(2k+1)^{3}$ equals the number of such cubes. Next we divide by the volume $|\Lambda|=(2k+3)^{3}\ell^{3}=(2k+3)^{3}m_{*}N/\rho$ and we take the limit $k\to\infty$ with all the other parameters fixed. For the Jellium term we use the convergence to the periodic energy as stated for instance in [Lew22, Lemma 32]. We obtain
@@
where $M_{\mathbb{Z}^{3}}=\lim_{x\to 0}(G_{1}(x)-|x|^{-1})$ is the Madelung constant of the lattice $\mathbb{Z}^{3}$. As a last step we take the limit $\ell\to\infty$, using that the periodic Jellium problem converges to $e_{\rm Jel}$ in (16), by [LLS19, Section V] and [Lew22, Theorem 40]. There is an additional factor $(\rho/m_{*})^{4/3}$ since the effective density of the periodic problem is $\rho/m_{*}$. Our final bound is
@@ 4. Proof of the lower bound in Theorem 2
The name “grand-canonical” comes from statistical physics and $\mu_{*}$ is interpreted as a “chemical potential”. For us it is an essential concept because the relaxation of the volume constraint will allow for the subsequent localization procedure (the volume constraint is global and not satisfied locally).
@@
In this first step we localize our problem to sets of finite size. For this we use the Graf–Schenker inequality [GS95, HLS09]. Locally we get the grand-canonical problem introduced in (28) because although we start with a globally neutral system ($|\Omega|=\rho|\Lambda|$), neutrality does not hold locally. The precise statement is the following.
@@
We show in the next lemma that with $C={\rm Per}\,(\Delta)$ we have
@@
Next we take a large cube $\Lambda_{L}:=(-L/2,L/2)^{3}$ and any domain $\Omega\subset\Lambda_{L}$. The Graf–Schenker inequality from [GS95] states that
@@
For this inequality to hold it is essential to average both over rotations and translations. For fixed $g$ we expand the integrand in the remaining double integral. Together with (30) and (31), we obtain for all $\Omega\subset\Lambda_{L}$
@@
The right side is the Jellium energy for particles with a charge $m_{*}$, in a fixed tetrahedron of volume $A^{3}$ and a uniform background of density 1. It is grand-canonical because we optimize over the number of particles $n$ instead of improsing the neutrality condition $m_{*}n=A^{3}$. It is however known that the optimal $n$ satisfies $n=A^{3}/m_{*}+o(A^{2})$. Fluctuations about the neutral case are negligible compared to the surface, otherwise the resulting energy is modified to the order of the volume (see [LL72, LN75, GS95a] and [Lew22, Cor. 50]).
We call (35) the “ultra-dilute” limit because we expect that the solution to the left side involves finitely many droplets (of the order of $A^{3}$) located at distances of order $\rho^{-1/3}$ in the very large domain $\ell\Delta$ of side length $\ell=A\rho^{-1/3}$. A similar limit was considered for the canonical problem in [CP10, KMN16, EFK20]. After rescaling everything by $\rho^{1/3}$ the droplets look like points and we find the right side.
@@
For $\rho=0$ we just recover the result from [FKN16], of which the lemma is an easy extension. When $\rho>0$ the inequality (36) says that $|\Omega|$ is at most of the same order as the background charge $\rho\,|\Lambda|\approx\rho\ {\rm diam}(\Lambda)^{3}$.
@@
Our proof follows the method in [FKN16]. For any $\nu\in\mathbb{S}^{2}$ and $t\in\mathbb{R}$, we define the half spaces and the plane
@@
Upon translating everything we can assume that $\Lambda\subset B(0,R)$ with $R=\text{diam}(\Lambda)$ (in fact, by Jung’s theorem [Fed96, Thm. 2.10.41] we could take $R=\text{diam}(\Lambda)\sqrt{3/8}$ which would slightly improve the constant $16\pi$ in the statement of the lemma). Then we know that $\mathcal{H}^{\pm}_{\nu,\pm t}\cap\Lambda=\emptyset$ for all $\pm t\geqslant R$. For every $t\in(-R,R)$, we split $\Omega$ into
@@
by Newton’s theorem as in (25). Next we integrate over $\nu\in\mathbb{S}^{2}$ and $t\in(-R,R)$, which gives as in [FKN16]
@@
Writing $A^{3}/m_{*}=N+t=(1-t)N+t(N+1)$ with $N\in\mathbb{N}_{0}$ and $0\leqslant t<1$ we can then bound the first minimum by the convex combination
@@
Next we remark as in [LD68, DL83] that the right side is linear with respect to each $m_{*}^{(j)}$, when all the other ones are fixed. We can thus replace all the $m_{*}^{(j)}$ by either 0 or the maximal value $m_{*}$ and decrease the right side. When they are replaced by 0 this is equivalent to discarding the corresponding points $y_{j}$, that is, to decrease $K$. Relabelling the points $y_{j}$ so that the $K^{\prime}$ first end up having mass $m_{*}$, we find that
@@
as we wanted. The grand-canonical Jellium problem has the same thermodynamic limit as its canonical counterpart in (16) by [Lew22, Thm. 24]. This concludes the proof of Theorem 2.∎
@@ Appendix A Proof of Theorem 1 on the thermodynamic limit
In this appendix we describe how to prove the existence of the thermodynamic limit stated in (6), using the method introduced by Lieb, Lebowitz and Narnhofer in [LL72, LN75]. It is also presented in [LS09, Chap. 14].
The proof is, as usual, divided into three steps. The first step consists of establishing the existence of the limit for a special sequence of balls $B_{j}$ growing exponentially fast. This is done by packing any $B_{K}$ by a certain disjoint union of copies of the smaller balls $B_{0},...,B_{K-1}$ in such a way that the remaining space has a volume negligible compared with the large volume $|B_{K}|$. The existence of such packings was proved first in [LL72] and has become known as a “swiss cheese theorem”. In the next two steps we use this special sequence to obtain the limit for any sequence of sets $\Lambda_{n}\nearrow\mathbb{R}^{3}$, satisfying suitable regularity conditions. Upper and lower bounds are proved using a similar swiss cheese for the inside or the outside of $\Lambda_{n}$.
@@ A.1. Construction of a quadrupole layer
We start by presenting a technical result that will be used in each of the three steps. It will allows us to deal with the neighborhood of the boundary of the sets we will manipulate. Recall that for the slowly-decaying Coulomb potential, a small charge disturbance can have a large effect far away. When constructing trial sets, it is therefore important to ensure that the charges locally balance well. Local neutrality is not enough and in the following proposition we will achieve an additional vanishing dipole condition. Handling this situation for the liquid drop model is harder than what was previously done in [LL72, LN75] and this is our main contribution in this appendix.
@@
Let $\rho\leqslant 1/2$. Let $\Lambda\subset\mathbb{R}^{3}$ be an (r,L)(r,L)–Lipschitz open set with compact boundary. Let $C_{z}=\varepsilon z+(-\varepsilon/2,\varepsilon/2)^{3}$ be a tiling of cubes of size $\varepsilon>0$, with $z\in\mathbb{Z}^{3}$. Define the outer cubic boundary layer of $\Lambda$ by
@@
Recall that an open set $\Lambda\subset\mathbb{R}^{3}$ is (r,L)(r,L)–Lipschitz if for any $x\in\partial\Lambda$, the set $\Lambda\cap B(x,r)$ is, in an appropriate coordinate system, equal to the epigraph of an LL–Lipschitz function. The interpretation of the condition (38) is that the charge distribution ${\mathds{1}}_{\Omega_{\alpha}}-\rho{\mathds{1}}_{\Lambda_{\alpha}}$ has no charge and no dipole moment. It can be thought as a small quadrupole, hence the title of the proposition. Recall that in our situation the quadrupole moment is the $3\times 3$ symmetric trace-less matrix
@@
Finally we come to the most interesting type of $\Lambda_{\alpha}$. We define each remaining $\Lambda_{\alpha}$ as the union of one $C_{z}$ with $z\in\mathcal{Z}_{i}$ together with all the small partial cubes $C^{\prime}_{z^{\prime}}\setminus\Lambda$ associated to it. The picture is therefore that we have a large full cube of volume $\varepsilon^{3}$ together with several small partial cubes of volume $\leqslant\varepsilon^{3}/K^{3}$ (Figure 2). The (r,L)(r,L)–Lipschitz property implies the following two facts, for $\varepsilon\leqslant r/C$:
@@
The constant $C$ only depends on $L$. The first claim (a) follows for instance from the fact that at any point $x\in\partial\Lambda$ there is a cone of height comparable with $r$ and opening angle bounded from below in terms of $L$, that is completely contained in $\mathbb{R}^{3}\setminus\Lambda$ (see for instance [Gri85, Thm. 1.2.2.2]). Hence for $\varepsilon\leqslant r/C$ this cone will contain a full cube $C_{z}$. For the second claim (b), we use that the $\eta$–thickened boundary inside a ball of radius $r^{\prime}$ satisfies